OpenStudy (brooke..help00):

Solve for x: 2 over 3 (x − 2) = 4x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (jackthegreatest):

\[\frac{ 2 }{ 3 (x-2) } = 4x\]

OpenStudy (jackthegreatest):

is that the equation?

OpenStudy (brooke..help00):

yes

OpenStudy (jackthegreatest):

\[\frac{ 2 }{ 3x-6 }=4x\]

OpenStudy (jackthegreatest):

\[2=4x(3x-6)\]

OpenStudy (brooke..help00):

ok:)

OpenStudy (jackthegreatest):

2=12x^2-24x

OpenStudy (jackthegreatest):

\[0=12x^2-24x-2\]

OpenStudy (jackthegreatest):

0=2(6x^2-12x-1)

OpenStudy (jackthegreatest):

are you able to factor that?

OpenStudy (jackthegreatest):

aare you sure the equation is right? because \[6x^2-12x-1 \] is not factorable

OpenStudy (brooke..help00):

\[\frac{ 2 }{ 3 }\left( x-2 \right)= 4x\]

OpenStudy (brooke..help00):

@jackthegreatest

OpenStudy (brooke..help00):

@OtherWorldly

OpenStudy (jackthegreatest):

ahhh see i knew something was wrong

OpenStudy (jackthegreatest):

\[2 (x-2) = 4x (3)\]

OpenStudy (jackthegreatest):

2x -4 =12x

OpenStudy (jackthegreatest):

-10x = 4

OpenStudy (jackthegreatest):

x = -2/5